6/x+7=8/x
What does x equal?

Answer by algebrahouse.com(1659)

6/x + 7 = 8/x
x[6/x + 7 = 8/x] {multiplied entire equation by x}
x(6/x) + x(7) = x(8/x) {multiplied each term by x}
6 + 7x = 8 {cancelled and multiplied what's left}
7x = 2 {subtracted 6 from both sides}
x = 2/7 {divided both sides by 7}

Unless it is supposed to be this:
6/(x + 7) = 8/x
8(x + 7) = 6x {two fractions equal to each other is a proportion, cross-multiplied}
8x + 56 = 6x {used distributive property}
56 = -2x {subtracted 8x from both sides}
x = -28 {divided both sides by -2}

For more help from me, visit: www.algebrahouse.com